A new business requires $100,000 investment. The project can either be entirely equity financed, or 70% of the investment can be funded by a bank loan at 10%. Assume no tax. EBIT is expected to be $25,000, but it could be as low as $3,000.

(a) calculate ROE for all 4 scenarios (unlevered low EBIT, unlevered high EBIT, levered low EBIT, unlevered high EBIT).
Answer #1

unlevered low EBIT

ROE = EBIT/Equity = 3000/100000 = 3%

unlevered high EBIT

ROE = EBIT/Equity = 25000/100000 = 25%

levered low EBIT

ROE = (EBIT-loan*interest)/(Equity before -loan) = (3000-100000*0.7*0.1)/(100000-100000*0.7)= -13.34%

levered high EBIT

ROE = (EBIT-loan*interest)/(Equity before -loan) = (25000-100000*0.7*0.1)/(100000-100000*0.7)= 60%